# 2cm2.5cm2.5cm2cm

### v2

@@ 3 Estimation and Asymptotic Theory
3.1 Estimators and Asymptotics
@@
Let $\boldsymbol{\theta}=(\theta_{1},\ldots,\theta_{6})^{\top}=(\beta_{0},\beta_{1},\beta_{2},\alpha_{0},\alpha_{1},\alpha_{2})^{\top}$ be the parameter vector, and $y_{1},\ldots,y_{n}$ be a realization of a NB tv-DINGARCH process $\{Y_{t}\}_{t=1}^{n}$. The conditional log-likelihood function of $Y_{2},\ldots,Y_{n}$ given $Y_{1}=y_{1}$ is given by $\ell(\boldsymbol{\theta})=\sum_{t=2}^{n}\ell_{t}(\boldsymbol{\theta})$, where
$\ell_{t}(\boldsymbol{\theta})=y_{t}[\log\lambda_{t}-\log(\lambda_{t}+\phi_{t})]+\phi_{t}[\log\phi_{t}-\log(\lambda_{t}+\phi_{t})]+\log\Gamma(y_{t}+\phi_{t})-\log\Gamma(\phi_{t})-\log y_{t}!,$
for $t=2,\ldots,n$, where we have omitted the dependence of $\lambda_{t}$ and $\phi_{t}$ on $\boldsymbol{\theta}$ for simplicity of notation. In practice, it is necessary to set some initial values for $\lambda_{1}$ and $\phi_{1}$, which are fixed in this paper. More specifically, we get such initial values based on the two first empirical moments of the count time series. This strategy has worked well as demonstrated in the simulated results to be discussed in the sequence. The conditional maximum likelihood estimator (CMLE) of $\boldsymbol{\theta}$, say $\widehat{\boldsymbol{\theta}}$, is given by $\widehat{\boldsymbol{\theta}}=\mbox{argmax}_{\boldsymbol{\theta}\in\Theta}\,\ell(\boldsymbol{\theta})$, where $\Theta=(0,\infty)\times[0,\infty)^{2}\times(0,\infty)\times[0,\infty)^{2}$ denotes the parameter space.
The score function associated to the conditional log-likelihood function is ${\bf U}(\boldsymbol{\theta})=\partial\ell(\boldsymbol{\theta})/\partial\boldsymbol{\theta}=\sum_{t=2}^{n}{\bf U}_{t}(\boldsymbol{\theta})$, with
@@
for $t=2,\ldots,n$, where $\Psi(x)=d\log\Gamma(x)/dx$, for $x>0$, is the digamma function. Explicit expressions for the derivatives involving $\lambda_{t}$ and $\phi_{t}$ are presented in the Supplementary Material. To establish the asymptotic normality of the CMLEs, the following proposition is necessary.
@@
Since $E(Y_{t}|\mathcal{F}_{t-1})=\lambda_{t}$ by definition, we immediately obtain that $E(S_{1t}(\boldsymbol{\theta})|\mathcal{F}_{t-1})=0$. We now compute the conditional expectation of $S_{2t}$ given $\mathcal{F}_{t-1}$, which involves $E\left(\Psi(Y_{t}+\phi_{t})|\mathcal{F}_{t-1}\right)$. For $a>0$ and $|c|<1$, we have that
@@
$E\left(\Psi(Y_{t}+\phi_{t})|\mathcal{F}_{t-1}\right)=\dfrac{(\lambda_{t}\phi_{t}^{-1}+1)^{-\phi_{t}}}{\Gamma(\phi_{t})}\sum_{y=0}^{\infty}\dfrac{\Gamma^{\prime}(y+\phi_{t})}{y!}\left(\dfrac{\lambda_{t}}{\lambda_{t}+\phi_{t}}\right)^{y}=\Psi(\phi_{t})-\log\left(\dfrac{\phi_{t}}{\lambda_{t}+\phi_{t}}\right),$
and then we obtain that $E(S_{2t}(\boldsymbol{\theta})|\mathcal{F}_{t-1})=0$. Hence, for all $t\geq 1$, $E({\bf U}_{t}(\boldsymbol{\theta})|\mathcal{F}_{t-1})=0$.
We will now show that ${\bf U}_{t}(\boldsymbol{\theta})$ is integrable. We have that $|S_{1t}|\leq|y_{t}/\lambda_{t}-1|\leq y_{t}/\beta_{0}+1$. Furthermore, $\partial\lambda_{t}/\partial\beta_{i}$ is a linear combination of $y_{t-1},\ldots,y_{1}$ and $\lambda_{1}$, for $i=0,1,2$. Using these results and under condition $\|{\bf A}\|_{1}+\|{\bf B}\|_{1}<1$, Theorem 2.2 gives that the score function associated with $\beta_{0},\beta_{1}$ and $\beta_{2}$ is integrable. We now study the score function associated with $\alpha_{0}$, $\alpha_{1}$ and $\alpha_{2}$. Note that $\left|-\dfrac{y_{t}-\lambda_{t}}{\lambda_{t}+\phi_{t}}\right|\leq\dfrac{y_{t}}{\lambda_{t}+\phi_{t}}+1\leq\dfrac{y_{t}}{2}\left(\dfrac{1}{\beta_{0}}+\dfrac{1}{\alpha_{0}}\right)+1$, where we have used that fact that $\dfrac{1}{a+b}\leq\dfrac{1}{2}\left(\dfrac{1}{a}+\dfrac{1}{b}\right)$ for $a,b>0$. Also, by using that $\dfrac{1}{2x}<\log x-\Psi(x)<\dfrac{1}{x}$ for $x>0$ (Alzer, 1997), we obtain that $0<\log\phi_{t}-\Psi(\phi_{t})<\phi_{t}^{-1}<\alpha_{0}^{-1}$. Similarly, it follows that $|\Psi(\lambda_{t}+\phi_{t})-\log(\lambda_{t}+\phi_{t})|<(\lambda_{t}+\phi_{t})^{-1}\leq\dfrac{1}{2}\left(\dfrac{1}{\beta_{0}}+\dfrac{1}{\alpha_{0}}\right)$. Therefore, $|S_{2t}|$ is bounded above by a linear combination of $y_{t}$. Moreover, $\dfrac{\partial\phi_{t}}{\partial\alpha_{j}}$ is a linear combination of $y_{t-1},\ldots,y_{1}$ and $\phi_{1}$, for $j=0,1,2$. Again, an application of Theorem 2.2 gives that the score function associated with $\alpha_{0}$, $\alpha_{1}$ and $\alpha_{2}$ is also integrable. ∎
Let us now discuss the asymptotic distribution of the CMLE for the NB tv-DINGARCH model. A more detailed technical result will be established in Theorem 3.1. By using Proposition 3.1 and the square integrability of ${\bf U}(\boldsymbol{\theta})$ (this can be demonstrated using very similar arguments to the proof of Proposition 3.1), we can apply the Central Limit Theorem for martingale difference (for instance, see Corollary 3.1 from Hall and Heyde (1980)) to obtain that
@@
as $n\rightarrow\infty$, with
${\bf J}_{1}(\boldsymbol{\theta})\equiv\dfrac{1}{n}\sum_{t=2}^{n}\mbox{Var}({\bf U}_{t}(\boldsymbol{\theta})|\mathcal{F}_{t-1})=\dfrac{1}{n}\sum_{t=2}^{n}\begin{pmatrix}b_{t}\dfrac{\partial\lambda_{t}}{\partial\boldsymbol{\theta}_{*}}\dfrac{\partial\lambda_{t}}{\partial\boldsymbol{\theta}_{*}}^{\top}&{\bf 0}\\ {\bf 0}&l_{t}\dfrac{\partial\phi_{t}}{\partial\boldsymbol{\theta}^{*}}\dfrac{\partial\phi_{t}}{\partial\boldsymbol{\theta}^{*}}^{\top}\end{pmatrix},$
$\boldsymbol{\theta}_{*}=(\theta_{1},\theta_{2},\theta_{3})^{\top}$, $\boldsymbol{\theta}^{*}=(\theta_{4},\theta_{5},\theta_{6})^{\top}$, $\Psi^{\prime}(x)=d\Psi(x)/dx$,
@@
Then, we apply the Law of Large Numbers for stationary and ergodic sequences to obtain that
@@
${\bf J}_{2}(\boldsymbol{\theta})\equiv\dfrac{1}{n}\sum_{t=2}^{n}E(-\nabla{\bf U}_{t}(\boldsymbol{\theta})|\mathcal{F}_{t-1})=\dfrac{1}{n}\sum_{t=2}^{n}\begin{pmatrix}b_{t}\dfrac{\partial\lambda_{t}}{\partial\boldsymbol{\theta}_{*}}\dfrac{\partial\lambda_{t}}{\partial\boldsymbol{\theta}_{*}}^{\top}&\bf 0\\ \bf 0&d_{t}\dfrac{\partial\phi_{t}}{\partial\boldsymbol{\theta}^{*}}\dfrac{\partial\phi_{t}}{\partial\boldsymbol{\theta}^{*}}^{\top}\end{pmatrix},$
@@
with $\boldsymbol{\Sigma}(\boldsymbol{\theta})=\boldsymbol{\Omega}_{2}^{-1}(\boldsymbol{\theta})\boldsymbol{\Omega}_{1}(\boldsymbol{\theta})\boldsymbol{\Omega}_{2}^{-1}(\boldsymbol{\theta})=\boldsymbol{\Omega}_{1}^{-1}(\boldsymbol{\theta})$. We have that $\widehat{\boldsymbol{\Sigma}}={\bf J}_{1}^{-1}(\widehat{\boldsymbol{\theta}})$ is a consistent estimator for $\boldsymbol{\Sigma}$. Another two consistent estimators for $\boldsymbol{\Sigma}$ are ${\bf S}^{-1}_{1}(\widehat{\boldsymbol{\theta}})$ and ${\bf S}^{-1}_{2}(\widehat{\boldsymbol{\theta}})$, where ${\bf S}_{1}(\boldsymbol{\theta})=n^{-1}\sum_{t=2}^{n}{\bf U}_{t}(\boldsymbol{\theta}){\bf U}_{t}(\boldsymbol{\theta})^{\top}$ and ${\bf S}_{2}(\boldsymbol{\theta})=-n^{-1}\sum_{t=2}^{n}\nabla{\bf U}_{t}(\boldsymbol{\theta})$.
There are required additional technical conditions to ensure the consistency and asymptotic normality of the CML estimators, which are provided in the next theorem, where its proof can be found in the Appendix. Before doing that, we have to consider lower and upper bounds for the possible values that the parameters can assume and then we define
=$\mathcal{B}=\{\boldsymbol{\theta}:\beta_{i,low}<\beta_{i}<\beta_{i,up},\,\,\alpha_{j,low}<\alpha_{j}<\alpha_{j,up},\,\,i,j=0,1,2\}.$ (8)
@@
Let $\{Y_{t}\}_{t=1}^{n}$ follow the linear negative binomial tv-DINGARCH(1,1,1,1) process. Assume that $\|{\bf A}\|_{1}+\|{\bf B}\|_{1}<1$ and that the true value of $\boldsymbol{\theta}$ is an interior point of $\mathcal{B}$. Then, there exists an open set $\mathcal{A}\subset\mathcal{B}$ such that the conditional log-likelihood function $\ell(\cdot)$ has a global maximum point on $\mathcal{A}$, say $\widehat{\boldsymbol{\theta}}$, with probability tending to 1 as $n\rightarrow\infty$. Furthermore, $\widehat{\boldsymbol{\theta}}$ is consistent for $\boldsymbol{\theta}$ and satisfy the asymptotic normality (7).

### v3

@@ 3 Estimation and Asymptotic Theory
Remark 3.1.
Some earlier work, in the context of negative-binomial quasi-likelihood inference for count time series models was reported by Aknouche et al. (2018). More precisely these authors studied model (3) but with a constant dispersion parameter. There are several differences between our work and that of Aknouche et al. (2018).
  1. 1.
    The first major difference is the type of models considered for the dispersion parameter. We assume a dynamic model–recall equation (3)–whereby the choice $\alpha_{1}=\alpha_{2}=0$ implies the model considered by Aknouche et al. (2018). So, in this sense, we extend this contribution.
  2. 2.
    The estimation method considered by Aknouche et al. (2018) is different from the direct maximization problem we are studying in this work. Indeed, Aknouche et al. (2018) consider a two-stage NB-QMLE (see their Sec. 3.2). We have developed theory for direct maximization of the negative binomial quasi-likelihood under an extended model.
  3. 3.
    In Aknouche et al. (2018) the invertibility of the Fisher information matrix is not proved but instead it is assumed; see Assumption (A9) (existence, finiteness, and invertibility of $J_{r}$ following the notation of their paper). The invertibility of the information matrix is addressed by our contribution, in a more general form, by the proof of Theorem 3.1.
  4. 4.
    Furthermore, these authors assume finiteness of moments (Aknouche et al., 2018, Assumption A3), which requires the existence of $\delta>1$ such that $E(X_{t}^{\delta})<\infty$. These are natural conditions to be assumed since the authors are studying exclusively quasi-likelihood inference. Instead we have proved such condition for this extended model we consider. There are additional differences, see for instance (Aknouche et al., 2018, Assumption A2) and compare with our Lemma A-1.
3.1 Likelihood Inference
@@
Denote by $\boldsymbol{\theta}_{1}=(\beta_{0},\beta_{1},\beta_{2})^{\top}$, $\boldsymbol{\theta}_{2}=(\alpha_{0},\alpha_{1},\alpha_{2})^{\top}$, and $\boldsymbol{\theta}=(\boldsymbol{\theta}_{1}^{\top},\boldsymbol{\theta}_{2}^{\top})^{\top}=(\theta_{1},\ldots,\theta_{6})^{\top}$ the parameter vector and suppose that $Y_{1},\ldots,Y_{n}$ is a sample from a NB tv-DINGARCH process $\{Y_{t}\}_{t=1}^{n}$. The conditional log-likelihood function of $Y_{2},\ldots,Y_{n}$ given $Y_{1}=y_{1}$ is given by (up to a constant)
$\ell(\boldsymbol{\theta})$ $=\sum_{t=2}^{n}\ell_{t}(\boldsymbol{\theta})$ (5)
$\ell_{t}(\boldsymbol{\theta})$ $=y_{t}[\log\lambda_{t}-\log(\lambda_{t}+\phi_{t})]+\phi_{t}[\log\phi_{t}-\log(\lambda_{t}+\phi_{t})]+\log\Gamma(y_{t}+\phi_{t})-\log\Gamma(\phi_{t}),$
for $t=2,\ldots,n$, where we have omitted the dependence of $\lambda_{t}$ and $\phi_{t}$ on $\boldsymbol{\theta}$ for simplicity.
The computation of the log-likelihood function (5) depends on the initial values of $\lambda_{0}$ and $\phi_{0}$. If we denote by $\tilde{\lambda}_{0}$ and $\tilde{\phi}_{0}$ some fixed or random starting values of the process (3), then the observed log-likelihood function is given by
~$\tilde{\ell}(\boldsymbol{\theta})$ $=\sum_{t=1}^{n}\tilde{\ell}_{t}(\boldsymbol{\theta})$ (6)
~$\tilde{\ell}_{t}(\boldsymbol{\theta})$ $=y_{t}[\log\tilde{\lambda}_{t}-\log(\tilde{\lambda}_{t}+\tilde{\phi_{t})}]+\tilde{\phi}_{t}[\log\tilde{\phi}_{t}-\log(\tilde{\lambda}_{t}+\tilde{\phi}_{t})]+\log\Gamma(y_{t}+\tilde{\phi}_{t})-\log\Gamma(\tilde{\phi}_{t}),$
with $\tilde{\lambda}_{t}=\beta_{0}+\beta_{1}Y_{t-1}+\beta_{2}\tilde{\lambda}_{t-1}$ and $\tilde{\phi}_{t}=\alpha_{0}+\alpha_{1}Y_{t-1}+\alpha_{2}\tilde{\phi}_{t-1}$.
The conditional maximum likelihood estimator (CMLE) of $\boldsymbol{\theta}$, say $\widehat{\boldsymbol{\theta}}$, is given by $\widehat{\boldsymbol{\theta}}=\mbox{argmax}_{\boldsymbol{\theta}\in\Theta}\tilde{\ell}(\boldsymbol{\theta})$, where $\Theta=(0,\infty)\times[0,\infty)^{2}\times(0,\infty)\times[0,\infty)^{2}$ denotes the parameter space.
In Lemma A-1 in the Appendix, we show that the log-likelihood function (6) is approximated by (5) with $(Y_{t},\lambda_{t},\phi_{t})_{t\in\mathbb{Z}}$ the stationary and ergodic process obtained by Theorem 2.1; see also Francq and Zakoïan (2004), Straumann and Mikosh (2006) and Christou and Fokianos (2014), among others. Note that (6) is the observed likelihood function but we will be employing (5) in what follows (including the proof of Theorem 3.1 about the asymptotic normality of the CMLEs) for simplicity of notation and without loss of generality. In practice, we obtain initial values by employing the two first empirical moments of the data. This strategy has worked well; see Sec. 3.2 for empirical evidence.
The score function associated with the conditional log-likelihood function is
${\bf U}(\boldsymbol{\theta})=\frac{\partial\ell(\boldsymbol{\theta})}{\partial\boldsymbol{\theta}}=\sum_{t=2}^{n}{\bf U}_{t}(\boldsymbol{\theta}),$ (7)
with
@@
for $t=2,\ldots,n$, where $\Psi(x)=d\log\Gamma(x)/dx$, for $x>0$, is the digamma function. Explicit expressions for the derivatives involving $\lambda_{t}$ and $\phi_{t}$ are presented in the Supplementary Material.
To establish the asymptotic normality of the CMLEs, the following lemma (its proof given in the Appendix) and proposition are necessary.
Lemma 3.1.
For $\|{\bf A}\|_{1}+\|{\bf B}\|_{1}<1$, the $t$-th term of the score function ${\bf U}_{t}(\boldsymbol{\theta})$ is integrable for all $t\geq 2$.
@@
We will first show that $E({\bf U}_{t}(\boldsymbol{\theta})|\mathcal{F}_{t-1})=0$. This fact and the integrability of ${\bf U}_{t}(\boldsymbol{\theta})$ stated in Lemma 3.1 will give us the desirable result. Since $E(Y_{t}|\mathcal{F}_{t-1})=\lambda_{t}$ by definition, we immediately obtain that $E(S_{1t}(\boldsymbol{\theta})|\mathcal{F}_{t-1})=0$. We compute the conditional expectation of $S_{2t}$ given $\mathcal{F}_{t-1}$, which involves the term $E\left(\Psi(Y_{t}+\phi_{t})|\mathcal{F}_{t-1}\right)$ and it is the most difficult to deal with. For $a>0$ and $|c|<1$, we have that
@@
$E\left(\Psi(Y_{t}+\phi_{t})|\mathcal{F}_{t-1}\right)=\dfrac{(\lambda_{t}\phi_{t}^{-1}+1)^{-\phi_{t}}}{\Gamma(\phi_{t})}\sum_{y=0}^{\infty}\dfrac{\Gamma^{\prime}(y+\phi_{t})}{y!}\left(\dfrac{\lambda_{t}}{\lambda_{t}+\phi_{t}}\right)^{y}=\Psi(\phi_{t})-\log\left(\dfrac{\phi_{t}}{\lambda_{t}+\phi_{t}}\right).$
Thus
$E(S_{2t}(\boldsymbol{\theta})|\mathcal{F}_{t-1})$ $=$ $E\left(-\dfrac{Y_{t}-\lambda_{t}}{\lambda_{t}+\phi_{t}}+\log\left(\dfrac{\phi_{t}}{\lambda_{t}+\phi_{t}}\right)+\Psi(y_{t}+\phi_{t})-\Psi(\phi_{t})\Big{|}\mathcal{F}_{t-1}\right)$
$=$ $-\dfrac{E(Y_{t}|\mathcal{F}_{t-1})-\lambda_{t}}{\lambda_{t}+\phi_{t}}+\log\left(\dfrac{\phi_{t}}{\lambda_{t}+\phi_{t}}\right)+E(\Psi(y_{t}+\phi_{t})|\mathcal{F}_{t-1})-\Psi(\phi_{t})=0.$
Hence, for all $t\geq 2$, $E({\bf U}_{t}(\boldsymbol{\theta})|\mathcal{F}_{t-1})=0$. From Lemma 3.1 and under the condition $\|{\bf A}\|_{1}+\|{\bf B}\|_{1}<1$, ${\bf U}_{t}(\boldsymbol{\theta})$ is integrable. Therefore the conclusion holds true. ∎
We next sketch the proof for obtaining the asymptotic distribution of the CMLE for the NB tv-DINGARCH model. A more detailed technical result will be established in Theorem 3.1. Recall (7). By using Proposition 3.1 and the square integrability of ${\bf U}(\boldsymbol{\theta})$ (this is proved using similar arguments to the proof of Proposition 3.1), we can apply the Central Limit Theorem for martingale difference (see Corollary 3.1 from Hall and Heyde (1980), for example) to obtain that
@@
as $n\rightarrow\infty$,
${\bf J}_{1}(\boldsymbol{\theta})\equiv\dfrac{1}{n}\sum_{t=2}^{n}\mbox{Var}({\bf U}_{t}(\boldsymbol{\theta})|\mathcal{F}_{t-1})=\dfrac{1}{n}\sum_{t=2}^{n}\begin{pmatrix}b_{t}\dfrac{\partial\lambda_{t}}{\partial{\boldsymbol{\theta}}_{1}}\dfrac{\partial\lambda_{t}}{\partial{\boldsymbol{\theta}}_{1}}^{\top}&{\bf 0}\\ {\bf 0}&l_{t}\dfrac{\partial\phi_{t}}{\partial{\boldsymbol{\theta}}_{2}}\dfrac{\partial\phi_{t}}{\partial{\boldsymbol{\theta}}_{2}}^{\top}\end{pmatrix},$
with
@@
Then, we apply the Law of Large Numbers for stationary and ergodic sequences (see Thm. 6.28 of Breiman (1992)) to obtain that
@@
${\bf J}_{2}(\boldsymbol{\theta})\equiv\dfrac{1}{n}\sum_{t=2}^{n}E(-\nabla{\bf U}_{t}(\boldsymbol{\theta})|\mathcal{F}_{t-1})=\dfrac{1}{n}\sum_{t=2}^{n}\begin{pmatrix}b_{t}\dfrac{\partial\lambda_{t}}{\partial{\boldsymbol{\theta}}_{1}}\dfrac{\partial\lambda_{t}}{\partial{\boldsymbol{\theta}}_{1}}^{\top}&{\bf 0}\\ {\bf 0}&d_{t}\dfrac{\partial\phi_{t}}{\partial{\boldsymbol{\theta}}_{2}}\dfrac{\partial\phi_{t}}{\partial{\boldsymbol{\theta}}_{2}}^{\top}\end{pmatrix},$
@@
with $\boldsymbol{\Sigma}(\boldsymbol{\theta})=\boldsymbol{\Omega}_{2}^{-1}(\boldsymbol{\theta})\boldsymbol{\Omega}_{1}(\boldsymbol{\theta})\boldsymbol{\Omega}_{2}^{-1}(\boldsymbol{\theta})=\boldsymbol{\Omega}_{1}^{-1}(\boldsymbol{\theta})$. We have that $\widehat{\boldsymbol{\Sigma}}={\bf J}_{1}^{-1}(\widehat{\boldsymbol{\theta}})$ is a consistent estimator for $\boldsymbol{\Sigma}$.
There are required additional technical conditions to ensure the consistency and asymptotic normality of the CML estimators, which are provided in the next theorem, whose proof can be found in the Appendix. We first consider lower and upper bounds for the possible values of the components of the parameter vector by introducing (recall that $\theta_{3}=\beta_{2}$ and $\theta_{6}=\alpha_{2}$)
=$\mathcal{B}=\Bigl{\{}\boldsymbol{\theta}:0<\theta_{i,low}\leq\theta_{i}\leq\theta_{i,up},i,=1,\ldots,6,\leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ \theta_{3,up}<1,\theta_{6,up}<1\Bigr{\}}.$ (11)
@@
Let $\{Y_{t}\}_{t=1}^{n}$ follow the linear negative binomial tv-DINGARCH(1,1,1,1) process. Assume that $\|{\bf A}\|_{1}+\|{\bf B}\|_{1}<1$ and that the true value of $\boldsymbol{\theta}$ is an interior point of $\mathcal{B}$. Then, there exists an open set $\mathcal{A}\subset\mathcal{B}$ such that the conditional log-likelihood function $\ell(\cdot)$ has a global maximum point on $\mathcal{A}$, say $\widehat{\boldsymbol{\theta}}$, with probability tending to 1 as $n\rightarrow\infty$. Furthermore, $\widehat{\boldsymbol{\theta}}$ is consistent for $\boldsymbol{\theta}$ and satisfies (10).